zelma is eighteen years older than her son. she was three times as old as he was one year ago. how old are they now?
Answer by richwmiller(17219)

Loved those age problems since I first saw them
z=18+s
z-1=3(s-1)
substitute
18+s-1=3(s-1)
17-s=3s-3
20=2s
s=10
z=28
check
28-1=3*(10-1)
27=3*9
27=27
ok